OpenStudy (anonymous):

A roller coaster travels 38.8 m at an angle of 14.0◦ above the horizontal. How far does it move horizontally? Answer in units of m

OpenStudy (anonymous):

This is a vectors problem. We are given magnitude (38.8) and direction (14.0°). We solve this by solving for the x-component of the vector. If x = x-component, y = y-component, M = magnitude, and D = Direction, \[x = M \cos D \] \[y = M \sin D\] So we plug in to the x one: \[x = M \cos D\] \[x = 38.8 \cos (14.0°)\] \[x \approx37.65\] The roller coaster traveled about 37.65 meters horizontally.

OpenStudy (anonymous):

thx